# Classical and quantum thermodynamics in a non–equilibrium regime: Application to Stirling engine

### v1

@@ III.4.1 Simulation details
Figure 2: Time profiles of (a) the electric field $x(t)$ and (b) the inverse temperature $\beta(t)$ for the Stirling engine. The red, green, blue, and orange lines represent (i) hot isothermal, (ii) isoelectric (from hot to cold), (iii) cold isothermal, and (iv) isoelectric (from cold to hot) processes, respectively. Time $t$ is normalized by the period of the cycle $\tau$.
We performed simulations using the anharmonic potential system employed in our previous study.Koyanagi and Tanimura (2024a) Thus, we considered a quartic anharmonic potential with the external interaction described as $X_{\rm A}(\hat{q})=\hat{q}$. The potential function in Eq. (32) is expressed as
@@
where the constants are given by $U_{2}=0.1$, $U_{3}=0.02$, and $U_{4}=0.05$. Numerical calculations were carried out to integrate the thermodynamic HEOM in the quantum cases and the thermodynamic Kramers equation in the classical cases. The detailed conditions for the numerical calculations, including the working parameters, are presented in Table 3 and in Ref. Koyanagi and Tanimura, 2024a
Table 3: Parameter values used for the simulations of the Stirling engine. Here, $dx$ and $dp$ are the mesh sizes for position and momentum, respectively, in Wigner space. The integers $N$ and $K$ are the cutoff numbers used in the LT-FPE.
@@
To set the trial functions defined in Eqs. (42) and (43), we chose $N_{\beta}=N_{x}=5$. Then, using the Nelder–Mead method, we minimized $\Xi_{\rm A}^{\rm target}[\beta^{\rm trial}(t),x^{\rm trial}(t)]$ in Eq. (III.3) with a cutoff time $\Delta t=1.0$. We evaluated $\Xi_{\rm A}^{\rm neq}(t)$ every 20 steps. The value of the quasi-static Planck potential $\Xi_{\rm A}^{\rm qst}(\beta^{\rm qst},\tilde{x}^{\rm qst})$ in Eq. (III.3) was obtained by integrating the thermodynamic LT-QFPE and the Kramers equation.
@@ III.4.2 Results
Figure 3: $x(t)$$X_{\rm A}^{\rm neq}(t)$ diagrams for the Stirling engine in the classical case (1, left column) and quantum case (2, right column) for (a) $A=0.5$ (weak), (b) $1.0$ (intermediate), and (c) $1.5$ (strong) SB coupling strengths. In each plot, the four curves (or lines) represent (i) hot isothermal (red), (ii) from hot to cold thermostatic (green), (iii) cold isothermal (blue), and (iv) from cold to hot thermostatic (orange) processes, respectively. The processes evolve in a counterclockwise fashion over time in a heat engine, whereas they evolve in a clockwise fashion over time in a refrigerator.
Figure 4: $T(t)$$S_{\rm A}^{\rm neq}(t)$ diagrams for the Stirling engine in the classical case (1, left column) and quantum case (2, right column) for (a) $A=0.5$ (weak), (b) $1.0$ (intermediate), and (c) $1.5$ (strong) SB coupling strength, respectively. Each cycle starts with the red arrow, and the four curves represent (i) hot isothermal (red), (ii) from hot to cold thermostatic (green), (iii) cold isothermal (blue), and (iv) from cold to hot thermostatic (orange) processes, respectively.
@@
The extensive variables $X_{\rm A}^{\rm neq}(t)$ and $U_{\rm A}^{\rm neq}(t)$ were obtained from Eqs. (37) and (38), respectively. The non–equilibrium entropy expressed as $S_{\rm A}^{\rm neq}(t)=k_{\rm B}\Lambda_{\rm A}^{\rm neq}(t)$ was then obtained from $X_{\rm A}^{\rm neq}(t)$, and $U_{\rm A}^{\rm neq}(t)$ and $\Xi_{\rm A}^{\rm neq}(t)$ were evaluated according to the procedure described in Sec. III.3 using the TDL transformations Eqs. (16) and (17). The results are depicted as non–equilibrium $x(t)$$X_{\rm A}^{\rm neq}(t)$ and $T(t)$$S_{\rm A}^{\rm neq}(t)$ diagrams, which are closed because the cycle is steady and because $T(t)$, $S_{\rm A}^{\rm neq}(t)$, $x(t)$, and $X_{\rm A}^{\rm neq}(t)$ are state variables.
@@
We first present the $x(t)$$X_{\rm A}^{\rm neq}(t)$ diagrams for weak ($A=0.5$), intermediate ($A=1.0$), and strong ($A=1.5$) system–bath coupling strengths in Fig. 3. The area enclosed by each diagram corresponds to positive work when evolving in the counterclockwise direction over time, whereas it corresponds to negative work when evolving clockwise. The work performed in one cycle is presented in Table 4. In each diagram in Fig. 3, the red and blue horizontal lines appear because there is a time delay in the change of $X_{\rm A}^{\rm neq}(t)$ with respect to $x(t)$ for the heat bath to take effect.
@@
As shown in Fig. 3 and Table 4, the larger the SB coupling, the smaller the quantum effect, because it is suppressed by relaxation.Ikeda and Tanimura (2019) Thus, when the coupling is large, dissipation becomes dominant, and the system becomes a refrigerator (or damper). This is because the time scales of fluctuations and dissipation are different: in the non–equilibrium case, dissipation dominates when the SB coupling is large, whereas in the quasi-static case, fluctuations and dissipation are balanced regardless of the coupling strength. This is a distinct difference from the Carnot case; the efficiency reaches a maximum in the intermediate SB coupling region where $A$ is neither large nor small.KT20JCP2
@@
Next, we present the $T(t)$$S_{\rm A}^{\rm neq}(t)$ diagrams in Fig. 4. The area enclosed by a clockwise curve represents the difference between extensive heat and waste heat, $Q_{\rm A}^{ext}-Q_{\rm A}^{\rm wst}$, per cycle. In the quasi-static case, the area agrees with the extensive heat per cycle because $Q^{\rm wst}$ becomes zero. In the thermostatic processes (green and orange curves), when the SB coupling is small, there is a time delay in the change of $S_{\rm A}^{\rm neq}(t)$ with respect to $T(t)$. This occurs because it takes time for the system to be excited when thermal fluctuations are small, whereas the delay is almost negligible when the SB coupling is strong.
@@
Figure 5: (a) Waste heat current $dQ^{\rm wst}(t)/dt$, (b) enthalpy, and (c) internal energy as functions of $t$. The blue, green, and red curves represent the weak ($A=0.5$), intermediate ($A=1.0$), and strong ($A=1.5$) SB coupling cases, respectively. The solid and dashed curves represent the quantum and classical results, respectively. The plots of waste heat currents have been time-averaged over 10 steps each using the raw data.
@@
In Fig. 5, we show the (time-averaged) waste heat current, enthalpy, and internal energy as functions of $t$. As we show in Sec. II, the waste heat current is always negative, and it becomes large in the isothermal case for larger SB coupling because of the strong dissipation, whereas it becomes small in the thermostatic processes for the case of weak SB coupling because fluctuations are suppressed.
@@
Using the graph of the waste heat current, we can improve the cycle efficiency. For instance, in the case of strong coupling, because a large waste heat current occurs in the isothermal processes, by increasing the duration of the isothermal processes, we can reduce the waste heat in the cycle.
@@
For weak SB coupling, both the enthalpy and internal energy in the classical case agree with those in the quantum case, whereas for strong SB coupling, they disagree, owing to the quantum coherence between the system and bath (bathentanglement).Tanimura (2020)
@@ IV Conclusions
The virtue of thermodynamics lies in its ability to describe, in a system-independent manner, macroscopic thermal phenomena resulting from complex microscopic interactions as changes in thermodynamic potentials described as interrelated intensive and extended variables through Legendre transformations.
We have developed the laws of thermodynamics as applied to work in a system-independent manner, based on the principle of DL minimum work. Subsequently, we have developed a non–equilibrium thermodynamic theory that describes non–equilibrium thermodynamic potential in terms of time-dependent intensive and extensive variables, along with waste heat.
Thus, the non–equilibrium Massieu–Planck and Helmholtz–Gibbs potentials are obtained in time-derivative form in terms of non–equilibrium extensive and intensive variables, which are state variables along with waste heat, which is not a state variable. Our results are summarized in Tables 1 and 2 and are consistent with traditional thermodynamics in a quasi-static case, as presented in Tables 5 and 6.
These results have been validated for quantum and classical Stirling engines through numerical simulations based on the thermodynamic SB model, which can describe both isothermal and thermostatic processes. An optimization algorithm has been used to evaluate non–equilibrium thermodynamic potentials. Work diagrams in non–equilibrium regimes have been presented and analyzed.
Although the time evolution of the system in the non–equilibrium regime is model-specific, it is possible to test our theory even in real systems through the use of optimization algorithms. By designing processes to reduce waste heat, it may be possible to develop efficient heat engines even for non–equilibrium processes.
In engineering, the effective energy from non-equilibrium to equilibrium is referred to as exergy.Rant (1956); Rosen and Dincer (2001) The non–equilibrium thermodynamic potentials we have introduced can be regarded as a generalization of it. The present theory provides a methodology to systematically evaluate and improve it, which should be useful from the perspective of Sustainable Development Goal (SDGs).
@@ Acknowledgments
The authors thank Yoshi Oono for his critical comments on the definition and terminology of thermodynamic variables, particularly for extensive variables. Y.T. was supported by JSPS KAKENHI (Grant No. B21H01884). S.K. was supported by JST fellowship, the establishment of university fellowships towards the creation of science technology innovation (Grant No. JPMJFS2123), and by Grant-in-Aid for JSPS Fellows (Grant No. 24KJ1373).
@@
The data that support the findings of this study are available from the corresponding author upon reasonable request.
@@ Appendix A Quasi-static thermodynamic potentials
Table 5: Total differential expressions for the quasi-static (qst.) entropic potentials as functions of the intensive variables $\beta^{\rm qst}(t)$ and $\tilde{x}^{\rm qst}(t)$ and the extensive variables $U_{\rm A}^{\rm qst}(t)$ and ${X}_{\rm A}^{\rm qst}(t)$. Entropy has two definitions, depending on whether the work variable is intensive or extensive. Of these DL entropies, the commonly used one, which we call Massieu entropy (M-entropy), involves only extensive variables and is denoted by $\Lambda_{\rm A}^{\rm qst}[U_{\rm A}^{\rm qst},{X}_{\rm A}^{\rm qst}]$, whereas the less widely used one, which we call Planck entropy (P-entropy), is denoted by $\Gamma_{\rm A}^{\rm qst}[U_{\rm A}^{\rm qst},\tilde{x}^{\rm qst}]$. Whereas the enthalpy $H_{\rm A}^{\rm qst}(t)$ was chosen as the natural variable in Ref. Koyanagi and Tanimura, 2024a, here we chose the internal energy $U_{\rm A}^{\rm qst}$ instead. Each potential is related to the others by the Legendre transformations shown in the final column.
@@
$\frac{d\tilde{Q}_{\rm A}^{int}(t)}{dt}\equiv\beta(t)\frac{dU_{\rm A}(t)}{dt}+X_{\rm A}(t)\frac{d\tilde{x}(t)}{dt}$ (45)
@@
The extensive and intensive work and heat $\tilde{W}_{\rm A}^{ext}(t)$, $\tilde{W}_{\rm A}^{int}(t)$, $\tilde{Q}_{\rm A}^{ext}(t)$, and $\tilde{Q}_{\rm A}^{int}(t)$ are interrelated through the Legendre transformations (6), (7), and (46). Therefore, the inequalities (8), (9),
@@
where $\Gamma_{\rm A}^{\rm qst}$ is the Planck entropy (P-entropy), are all expressions of the second law of thermodynamics.
The total differential forms of the Massieu–Planck potentials are presented in Table 5.
The Helmholtz–Gibbs potentials can be obtained form the Massieu–Planck potentials using the definitions $F^{\rm qst}_{\rm A}(t)=-\Phi_{\rm A}^{\rm qst}(t)/\beta^{\rm qst}(t)$ and $G^{\rm qst}_{\rm A}(t)=-\Xi_{\rm A}^{\rm qst}(t)/\beta^{\rm qst}(t)$. From Eq. (6), we have
@@
Table 6: Total differential expressions for the quasi-static (qst.) thermodynamic potentials as functions of the intensive variables $T^{\rm qst}(t)$ and $x^{\rm qst}(t)$ and the extensive variables $S_{\rm A}^{\rm qst}(t)$ and ${X}_{\rm A}^{\rm qst}(t)$, which are related through the Legendre transformations shown in the final column.Koyanagi and Tanimura (2024a)
Helmholtz $dF_{\rm A}^{\rm qst}=-S_{\rm A}^{\rm qst}dT^{\rm qst}+x^{\rm qst}dX_{\rm A}^{\rm qst}$ $T^{\rm qst},X_{\rm A}^{\rm qst}$cdots
Gibbs $dG_{\rm A}^{\rm qst}=-S_{\rm A}^{\rm qst}dT^{\rm qst}-X_{\rm A}^{\rm qst}dx^{\rm qst}$ $T^{\rm qst},x^{\rm qst}$ $G_{\rm A}^{\rm qst}=F_{\rm A}^{\rm qst}-x^{\rm qst}X_{\rm A}^{\rm qst}$
Internal $dU_{\rm A}^{\rm qst}=T^{\rm qst}dS_{\rm A}^{\rm qst}+x^{\rm qst}dX_{\rm A}^{\rm qst}$ $S_{\rm A}^{\rm qst},X_{\rm A}^{\rm qst}$ $U_{\rm A}^{\rm qst}=F_{\rm A}^{\rm qst}+T^{\rm qst}S_{\rm A}^{\rm qst}$
Enthalpy $dH_{\rm A}^{\rm qst}=T^{\rm qst}dS_{\rm A}^{\rm qst}-X_{\rm A}^{\rm qst}dx^{\rm qst}$ $S_{\rm A}^{\rm qst},x^{\rm qst}$ $H^{\rm qst}_{\rm A}=G_{\rm A}^{\rm qst}+T^{\rm qst}S_{\rm A}^{\rm qst}$
Appendix B DL non–equilibrium minimum work principle

### v2

@@ III.4.1 Simulation details
Figure 2: Time profiles of (a) the electric field $x(t)$ and (b) the inverse temperature $\beta(t)$ for the thermostatic Stirling engine. The red, green, blue, and orange lines represent (i) hot isothermal, (ii) isoelectric ( hot to cold), (iii) cold isothermal, and (iv) isoelectric (cold to hot) processes, respectively.Koyanagi and Tanimura (2024c) The time $t$ is normalized by the period of the cycle $\tau$.
We performed simulations using the anharmonic potential system employed in our previous studies.Koyanagi and Tanimura (2024b, c) Thus, we considered a quartic anharmonic potential with the external interaction described as $X_{\rm A}(\hat{q})=\hat{q}$. The potential function in Eq. (32) is expressed as
@@
where the constants are given by $U_{2}=0.1$, $U_{3}=0.02$, and $U_{4}=0.05$. Numerical calculations were performed to integrate the T-QFPE in the quantum cases and the T-KE in the classical cases. The detailed conditions for the numerical calculations, including the working parameters, are presented in Table 3 and in Ref. Koyanagi and Tanimura, 2024b. Source code and results for the quasi-static case are presented in Ref. Koyanagi and Tanimura, 2024c.
Table 3: Parameter values used for the simulations of the thermostatic Stirling engine. Here, $dx$ and $dp$ are the mesh sizes for position and momentum, respectively, in Wigner space. The integers $N$ and $K$ are the cutoff numbers used in the T-QFPE.
@@
To set the trial functions defined in Eqs. (42) and (43), we chose $N_{\beta}=N_{x}=5$. Then, using the Nelder–Mead method, we minimized $\Xi_{\rm A}^{\rm target}[\beta^{\rm trial}(t),x^{\rm trial}(t)]$ in Eq. (III.3) with a cutoff time $\Delta t=1.0$. We evaluated $\Xi_{\rm A}^{\rm neq}(t)$ in 20 steps. The value of the quasi-static Planck potential $\Xi_{\rm A}^{\rm qst}(\beta^{\rm qst},\tilde{x}^{\rm qst})$ in Eq. (III.3) was obtained by integrating the thermodynamic T-QFPE and the T-KE.
@@ III.4.2 Results
Figure 3: $x(t)$$X_{\rm A}^{\rm neq}(t)$ diagrams for the thermostatic Stirling engine in the classical case (1, left column) and quantum case (2, right column) for (a) $A=0.5$ (weak), (b) $1.0$ (intermediate), and (c) $1.5$ (strong) SB coupling strengths. In each plot, the four curves (or lines) represent (i) hot isothermal (red), (ii) hot to cold thermostatic (green), (iii) cold isothermal (blue), and (iv) cold to hot thermostatic (orange) processes, respectively. These processes evolve in a counterclockwise manner over time in a heat engine, whereas they evolve in a clockwise manner over time in a refrigerator.
Figure 4: $T(t)$$S_{\rm A}^{\rm neq}(t)$ diagrams for the thermostatic Stirling engine in the classical case (1, left column) and quantum case (2, right column) for (a) $A=0.5$ (weak), (b) $1.0$ (intermediate), and (c) $1.5$ (strong) SB coupling strengths, respectively. Each cycle starts with a red arrow, and the four curves represent (i) hot isothermal (red), (ii) from hot to cold thermostatic (green), (iii) cold isothermal (blue), and (iv) from cold to hot thermostatic (orange) processes, respectively.
@@
Extensive variables $X_{\rm A}^{\rm neq}(t)$ and $U_{\rm A}^{\rm neq}(t)$ were obtained from Eqs. (37) and (38), respectively. The non-equilibrium entropy expressed as $S_{\rm A}^{\rm neq}(t)=k_{\rm B}\Lambda_{\rm A}^{\rm neq}(t)$ was then obtained from $X_{\rm A}^{\rm neq}(t)$, and $U_{\rm A}^{\rm neq}(t)$ and $\Xi_{\rm A}^{\rm neq}(t)$ were evaluated according to the procedure described in Sec. III.3 using the TDL transformations given by Eqs. (16) and (17). The results are depicted as non-equilibrium $x(t)$$X_{\rm A}^{\rm neq}(t)$ and $T(t)$$S_{\rm A}^{\rm neq}(t)$ diagrams, whose cycle trajectories are closed because the cycle is stationary and because $T(t)$, $S_{\rm A}^{\rm neq}(t)$, $x(t)$, and $X_{\rm A}^{\rm neq}(t)$ are state variables.
We first present the $x(t)$$X_{\rm A}^{\rm neq}(t)$ diagrams for weak ($A=0.5$), intermediate ($A=1.0$), and strong ($A=1.5$) SB coupling strengths in Fig. 3. The results for the quasi-static case are presented in Fig. 5 in Ref. Koyanagi and Tanimura, 2024c.
@@
The area enclosed by each diagram corresponds to positive work when evolving in the counterclockwise direction, whereas negative work corresponds to clockwise evolution. The work performed in one cycle is presented in Table 4. In each diagram in Fig. 3, the red and blue horizontal lines appear because a time delay exists in the change of $X_{\rm A}^{\rm neq}(t)$ with respect to $x(t)$ for the heat bath to take effect.
@@
As shown in Fig. 3 and Table 4, the larger the SB coupling, the smaller the quantum effect because it is suppressed by relaxation.Ikeda and Tanimura (2019) Thus, when the coupling is large, dissipation dominates, and the system becomes a refrigerator (or damper). This is because the time scales of fluctuation and dissipation are differ: in the non-equilibrium case, dissipation dominates when SB coupling is large, whereas in the quasi-static case (Fig. 5 in Ref. Koyanagi and Tanimura, 2024c), fluctuations and dissipation are balanced regardless of the coupling strength. This is a distinct difference from the Carnot case; the efficiency reaches a maximum in the intermediate SB coupling region, where $A$ is neither large nor small.Koyanagi and Tanimura (2022b)
@@
Next, we present the $T(t)$$S_{\rm A}^{\rm neq}(t)$ diagrams in Fig. 4. The results for the quasi-static case are presented in Fig. 6 in Ref. Koyanagi and Tanimura, 2024c. The area enclosed by the clockwise curve represents the difference between the extensive heat and waste heat, $Q_{\rm A}^{ext}-Q_{\rm A}^{\rm wst}$, per cycle. In the quasi-static case (Fig. 6 in Ref. Koyanagi and Tanimura, 2024c), the area agree with the extensive heat per cycle because $Q^{\rm wst}$ becomes zero. In the thermostatic processes (green and orange curves), when SB coupling is small, a time delay is observed in the change in $S_{\rm A}^{\rm neq}(t)$ with respect to $T(t)$. This occurs because it takes time for the system to be excited when thermal fluctuations are small, whereas the delay is almost negligible when SB coupling is strong.
@@
Figure 5: (a) Entropy production rate $d\Sigma_{\rm A}(t)/dt$, (b) the change of enthalpy $\Delta H_{\rm A}(t)$, and (c) internal energy $\Delta U_{\rm A}(t)$, as functions of $t$. The blue, green, and red curves represent the weak ($A=0.5$), intermediate ($A=1.0$), and strong ($A=1.5$) SB coupling cases, respectively. The solid and dashed curves represent the quantum and classical results, respectively. The plots of the entropy production rate were time-averaged over 21 steps each using the raw data.
@@
In Fig. 5, we show the (time-averaged) entropy production rate, enthalpy, and internal energy as functions of $t$. As we show in Sec. II, the entropy production rate is always positive, and becomes large in the isothermal case for larger SB coupling because of strong dissipation, whereas it becomes small in the thermostatic processes for the case of weak SB coupling because fluctuations are suppressed.
@@
Using the graph of the entropy production rate, we can improve the cycle efficiency. For instance, in the case of strong coupling, because a large entropy production rate occurs in isothermal processes, increasing the duration of the isothermal processes reduces the entropy production in the cycle.
@@
For weak SB coupling, both the enthalpy and internal energy in the classical case agree with those in the quantum case, whereas for strong SB coupling, they disagree because of bathentanglement.Tanimura (2020)
@@ IV Conclusions
The virtue of thermodynamics lies in its ability to describe, in a system-independent manner, macroscopic thermal phenomena resulting from complex microscopic interactions as changes in thermodynamic potentials that are described as interrelated intensive and extensive variables via Legendre transformations.
We developed the laws of thermodynamics as applied to work in a system-independent manner, based on the principle of DL minimum work. Subsequently, we have developed a non-equilibrium thermodynamic theory that describes the non-equilibrium Massieu–Planck potentials and Helmholtz–Gibbs potentials in time derivative form in terms of non-equilibrium extensive and intensive variables, which are state variables along with entropy production, which is not a state variable. Our results are summarized in Tables 1 and 2 and are consistent with traditional thermodynamics in a quasi-static case, as presented in Tables 5 and 6.
These results have been validated for thermostatic quantum and classical Stirling engines through numerical simulations based on the thermodynamic SB model, which can describe both isothermal and thermostatic processes. An optimization algorithm was used to evaluate non-equilibrium thermodynamic potentials. Work diagrams in non-equilibrium regimes were presented and analyzed.
In this paper, to perform numerical calculations, the external field was chosen as the intensive variable and the polarization as the extensive variable. In a Brownian model defined by a potential function, it is also possible to select pressure and volume as intensive and extensive variables, as the partition function was evaluated for an ideal gas.
Although our theory in the non-equilibrium regime is model-specific, it is possible to test it even in real systems by using optimization algorithms. By designing processes to reduce entropy production, efficient heat engines can be developed for non-equilibrium processes.
In engineering, the effective energy from non-equilibrium to equilibrium is referred to as exergy.Z The non-equilibrium thermodynamic potentials introduced can be regarded as a generalization of this concept. This theory provides a methodology to systematically evaluate and improve it, which should be useful from the perspective of Sustainable Development Goals (SDGs).
@@ Acknowledgments
The authors thank Yoshi Oono for his critical comments on the definition and terminology of thermodynamic variables, particularly extensive variables. Y.T. was supported by JSPS KAKENHI (Grant No. B21H01884). S.K. was supported by a JST fellowship, the establishment of university fellowships towards the creation of science technology innovation (Grant No. JPMJFS2123), and by Grant-in-Aid for JSPS Fellows (Grant No. 24KJ1373).
@@
Author Contributions
Shoki Koyanagi: Conceptualization (supporting); Formal analysis (lead); Investigation (equal); Software (lead); Writing - original draft (equal). Yoshitaka Tanimura: Conceptualization (lead); Formal analysis (supporting); Investigation (equal); Writing - review and editing (lead).
@@
The data that support the findings of this study are available from the corresponding authors upon reasonable request.
@@ Appendix A Quasi-static thermodynamic potentials
Table 5: Total differential expressions for the quasi-static (qst.) entropic potentials as functions of the intensive variables $\beta^{\rm qst}(t)$ and $\tilde{x}^{\rm qst}(t)$ and the extensive variables $U_{\rm A}^{\rm qst}(t)$ and ${X}_{\rm A}^{\rm qst}(t)$. Entropy has two definitions, depending on whether the work variable is intensive or extensive. Of these DL entropies, the commonly used one, which we call Massieu entropy (M-entropy), involves only extensive variables and is denoted by $\Lambda_{\rm A}^{\rm qst}[U_{\rm A}^{\rm qst},{X}_{\rm A}^{\rm qst}]$, whereas the less widely used one, which we call Planck entropy (P-entropy), is denoted by $\Gamma_{\rm A}^{\rm qst}[U_{\rm A}^{\rm qst},\tilde{x}^{\rm qst}]$. Whereas the enthalpy $H_{\rm A}^{\rm qst}(t)$ was chosen as the natural variable in Ref. Koyanagi and Tanimura, 2024b, here we chose the internal energy $U_{\rm A}^{\rm qst}$ instead. Each potential is related to the others by the Legendre transformations shown in the final column.
@@
$\frac{d\tilde{Q}_{\rm A}^{int}(t)}{dt}\equiv\beta(t)\frac{dU_{\rm A}(t)}{dt}+X_{\rm A}(t)\frac{d\tilde{x}(t)}{dt},$ (45)
@@
Extensive and intensive work and heat $\tilde{W}_{\rm A}^{ext}(t)$, $\tilde{W}_{\rm A}^{int}(t)$, $\tilde{Q}_{\rm A}^{ext}(t)$, and $\tilde{Q}_{\rm A}^{int}(t)$ are interrelated via Legendre transformations (6), (7), and (46). Therefore, we obtain the inequalities (8), (9),
@@
where $\Gamma_{\rm A}^{\rm qst}$ is the Planck entropy (P-entropy). These are all expressions of the second law of thermodynamics. The total differential forms of the Massieu–Planck potentials are presented in Table 5.
The Helmholtz–Gibbs potentials can be obtained from the Massieu–Planck potentials using the definitions $F^{\rm qst}_{\rm A}(t)=-\Phi_{\rm A}^{\rm qst}(t)/\beta^{\rm qst}(t)$ and $G^{\rm qst}_{\rm A}(t)=-\Xi_{\rm A}^{\rm qst}(t)/\beta^{\rm qst}(t)$. From Eq. (6), we obtain
@@
Table 6: Total differential expressions for the quasi-static (qst.) thermodynamic potentials as functions of intensive variables $T^{\rm qst}(t)$ and $x^{\rm qst}(t)$ and extensive variables $S_{\rm A}^{\rm qst}(t)$ and ${X}_{\rm A}^{\rm qst}(t)$. The potentials are related through the Legendre transformations shown in the final column
.Koyanagi and Tanimura (2024b) Qst. potential Differential form Natural var. Legendre transformation Helmholtz $dF_{\rm A}^{\rm qst}=-S_{\rm A}^{\rm qst}dT^{\rm qst}+x^{\rm qst}dX_{\rm A}^{\rm qst}$ $T^{\rm qst},X_{\rm A}^{\rm qst}$cdots Gibbs $dG_{\rm A}^{\rm qst}=-S_{\rm A}^{\rm qst}dT^{\rm qst}-X_{\rm A}^{\rm qst}dx^{\rm qst}$ $T^{\rm qst},x^{\rm qst}$ $G_{\rm A}^{\rm qst}=F_{\rm A}^{\rm qst}-x^{\rm qst}X_{\rm A}^{\rm qst}$ Internal $dU_{\rm A}^{\rm qst}=T^{\rm qst}dS_{\rm A}^{\rm qst}+x^{\rm qst}dX_{\rm A}^{\rm qst}$ $S_{\rm A}^{\rm qst},X_{\rm A}^{\rm qst}$ $U_{\rm A}^{\rm qst}=F_{\rm A}^{\rm qst}+T^{\rm qst}S_{\rm A}^{\rm qst}$ Enthalpy $dH_{\rm A}^{\rm qst}=T^{\rm qst}dS_{\rm A}^{\rm qst}-X_{\rm A}^{\rm qst}dx^{\rm qst}$ $S_{\rm A}^{\rm qst},x^{\rm qst}$ $H^{\rm qst}_{\rm A}=G_{\rm A}^{\rm qst}+T^{\rm qst}S_{\rm A}^{\rm qst}$
Appendix B DL non-equilibrium minimum work principle